# VLind-Bench: Measuring Language Priors in Large Vision-Language Models

### v3

@@ 5.1 Metrics
Here, $i$ is the data index, $N$ is the number of total instances in our benchmark, and $M_{\mathcal{T}}$ is the number of instances that have passed all the previous tests before $\mathcal{T}$ (which is essentially the number of instances considered by $\mathcal{T}$). To be more concise, $M_{\textrm{CK}}=M_{\textrm{VP}}=N$, $M_{\textrm{CB}}=|\{i\mid P_{\textrm{CK}}^{i}=1\}|$ and $M_{\textrm{LP}}=|\{i\mid P_{\textrm{CB}}^{i}=1\land P_{\textrm{VP}}^{i}=1\}|$. We refer to these four scores as pipeline scores, as they reflect the pipelined evaluation structure of VLind-Bench (columns under “Pipeline Score” in Table 2). Alternatively, following the commonly accepted definition of accuracy, the performance can be expressed as the ratio of correct instances to the total number of instances (columns under “Accuracy” in Table 2).
@@ 5.3 Main Results
The overall model performance is shown in Table 2. Surprisingly, numerous models demonstrated somewhat low scores in $S_{\textrm{CK}}$, implying a deficiency of commonsense knowledge in LVLMs. Conversely, $S_{\textrm{VP}}$ scores concerning object recognition ability exhibited relatively high scores. This pattern of low commonsense knowledge scores and high visual perception scores aligns with observations from previous work [31]. Additionally, the lower $S_{\textrm{CB}}$ and CB scores compared to $S_{\textrm{CK}}$ indicate that LVLMs are reluctant to respond contrary to commonsense knowledge.
@@
Finally, the generally low $S_{\textrm{LP}}$ score suggests that all models, except for GPT-4o, exhibit a reliance on language priors. This reliance was more pronounced in open-source models compared to proprietary ones. Furthermore, the reliance on language priors appeared inversely proportional to the scale of the backbone LLM. This trend can be observed by comparing the $S_{\textrm{LP}}$ scores across various sizes of models within the same LLaVA and InstructBLIP series.
RLHF-V
An exception to such trend between model scale and language prior is the superior performance of models that applied the RLHF-V [28] methodologies. Models such as OmniLMM and MiniCPM trained using RLHF-V, demonstrated superior performance compared to models of similar or greater scale. Specifically, RLHF-V employs a method called Dense Direct Preference Optimization (DDPO) to mitigate multimodal hallucination. DDPO constructs win-lose pairs by having humans modify only the hallucinatory spans in the model responses to align with image information, thereby forcing the use of visual modality to increase the reward. Such construction of training data might be the reason for the reduced reliance on language prior. Additionally, the high performance of these methods on counterfactual images suggests that the ability to utilize image information generalizes to out-of-distribution samples. Applying RLAIF-V [29], an AI-feedback variant of RLHF-V, to LLaVA 1.5 7B also results in significant performance improvement.
Table 3: Performance of selected models for different concepts.
@@
6 Discussion and Conclusion
@@
The tendency for the reliance on language prior to be inversely proportional to the scale of backbone LLMs may appear counterintuitive (i.e., LLaVA in Table 2). We have not identified the precise cause of this trend. One possible explanation is that larger pre-trained models are less prone to overfitting to the dataset during the visual instruction tuning process, thereby better maintaining their ability to attend to image information. In the experiments, we employ models with various scales of image encoders (ranging from approximately 300M to 5B), however, no clear correlation was observed between the language prior and the size of the image encoder.
@@
Limitations and Future Work
Although VLind-Bench minimized potential confounding factors in assessing language priors, there may still be unconsidered factors. The text and image data are sampled from generative models, which may result in discrepancies from real-world distributions. When evaluating LVLMs, we prompted the models to respond exclusively with either “True” or “False.” However, some models could perform better by generating a rationale before responding [25]. This aspect was not explored in our study but may be considered in future research.
Additionally, the CBs in our benchmark does not necessarily need to receive both text and image as input to check the commonsense bias. Such design choice is mostly due to a lack of established practices for feeding text-only inputs to LVLMs. As alternatives to $I_{\textrm{cf}}$, we conducted experiments using a plain single-color image or rendered text prompts as visual input (refer to Appendix C); however, none of these approaches works – these kinds of images can be considered out-of-distribution samples, and some proprietary models output error messages for these inputs. Exploring more established methods for text-only inputs in LVLMs falls outside the scope of our paper, but further research in this area is necessary both from a practical perspective and for a deeper understanding of how individual components of LVLMs operate.
Finally, while our primary goal in Section 4 was to generate data for a benchmarking purpose, we can also use this process to generate training data automatically. Training LVLMs with such dataset could help mitigate reliance on language priors, but we leave this as future work.
@@ Appendix A Human Verification and Model Prompt Details
Appendix B Data Samples
Appendix C Experiments Using a Plain White Image and Rendered Text Prompts
@@
LLaVA-NEXT 72B (Qwen 1.5 72B Chat) 94.4 95.7 76.1 58.6 75.5 46.7
OmniLMM 12B (Zephyr 7B $\beta$) 88.1 97.7 78.6 81.4 79.5 66.4
@@
LLaVA-NEXT 72B (Qwen 1.5 72B Chat) 88.4 95.7 74.9 54.8 74.2 46.7
OmniLMM 12B (Zephyr 7B $\beta$) 79.1 97.7 72.4 81.0 72.8 66.4
@@
LLaVA-NEXT 72B (Qwen 1.5 72B Chat) 89.1 95.7 70.6 54.2 72.8 46.7
OmniLMM 12B (Zephyr 7B $\beta$) 74.2 97.7 65.2 77.4 68.2 66.4
@@
Appendix D Model Performance by Image Style
@@
LLaVA-NEXT 72B (Qwen 1.5 72B Chat) 94.6 95.8 77.2 65.0 76.5 52.4
OmniLMM 12B (Zephyr 7B $\beta$) 88.8 97.7 81.8 82.8 83.1 70.5
@@
LLaVA-NEXT 72B (Qwen 1.5 72B Chat) 94.3 96.2 78.5 59.1 77.8 47.3
OmniLMM 12B (Zephyr 7B $\beta$) 88.5 98.1 81.4 80.4 82.4 67.7
@@
LLaVA-NEXT 72B (Qwen 1.5 72B Chat) 94.8 95.5 78.8 58.2 78.8 48.0
OmniLMM 12B (Zephyr 7B $\beta$) 87.7 97.8 82.2 82.0 82.5 68.4
Here, we observed how performance varies across different image styles. As mentioned in Section 4, we generated images in photorealistic, illustration, and cartoon styles. Table 5 shows that the performance across these styles in the CK, VP, and CB did not vary significantly. A notable variation in performance was observed only in LP, where the photorealistic style yielded better results compared to the other two styles. This could be due to the model’s assessment that images in the illustration or cartoon styles lack realism compared to photorealistic images, leading it to generate responses that align more closely with common sense.
Appendix E Data Access and License
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Dataset DOI: 10.57967/hf/2475

### v4

@@ 5.1 Metrics
Here, $i$ is the data index, $N$ is the number of total instances in our benchmark, and $M_{\mathcal{T}}$ is the number of instances that have passed all the previous tests before $\mathcal{T}$ (which is essentially the number of instances considered by $\mathcal{T}$). To be more concise, $M_{\textrm{CK}}=M_{\textrm{VP}}=N$, $M_{\textrm{CB}}=|\{i\mid P_{\textrm{CK}}^{i}=1\}|$ and $M_{\textrm{LP}}=|\{i\mid P_{\textrm{CB}}^{i}=1\land P_{\textrm{VP}}^{i}=1\}|$. We refer to these four scores as pipeline scores, as they reflect the pipelined evaluation structure of VLind-Bench (columns under “Pipeline Score” in Table 3). Alternatively, following the common definition of accuracy, the performance can be expressed as the ratio of correct instances to the total number of instances (columns under “Accuracy” in Table 3).
@@ 5.3 Main Results
The overall model performance is shown in Table 3. Surprisingly, numerous models demonstrated somewhat low scores in $S_{\textrm{CK}}$, implying a deficiency of commonsense knowledge in LVLMs. Conversely, $S_{\textrm{VP}}$ scores concerning object recognition ability exhibited relatively high scores. This pattern of low commonsense knowledge scores and high visual perception scores aligns with observations from previous work (Zhou et al., 2023). Additionally, the lower $S_{\textrm{CB}}$ and CB scores compared to $S_{\textrm{CK}}$ indicate that LVLMs are reluctant to respond contrary to commonsense knowledge.
@@
Finally, the generally low $S_{\textrm{LP}}$ score suggests that all models, except for GPT-4o, exhibit a reliance on language priors. This reliance was more pronounced in open-source models compared to proprietary ones. The reliance on language priors appeared inversely proportional to the scale of the backbone LLM. This trend can be observed by comparing the $S_{\textrm{LP}}$ scores across various sizes of models within the same LLaVA and InstructBLIP series.
To verify the validity of the VLind-Bench, we conducted experiments on a small handcrafted evaluation set, and the results are provided in Appendix D.
@@
Table 4: Performance of selected models for different concepts.
RLHF-V
An exception to such trend between model scale and language prior is the superior performance of models that applied the RLHF-V (Yu et al., 2024a) methodologies. Models such as OmniLMM and MiniCPM trained using RLHF-V, demonstrated superior performance compared to models of similar or greater scale. Specifically, RLHF-V employs a method called Dense Direct Preference Optimization (DDPO) to mitigate multimodal hallucination. DDPO constructs win-lose pairs by having humans modify only the hallucinatory spans in the model responses to align with image information, thereby forcing the use of visual modality to increase the reward. Such construction of training data might be the reason for the reduced reliance on language prior. Additionally, the high performance of these methods on counterfactual images suggests that the ability to utilize image information generalizes to out-of-distribution samples. Applying RLAIF-V (Yu et al., 2024b), an AI-feedback variant of RLHF-V, to LLaVA 1.5 7B also results in significant performance improvement.
@@ 6 Discussion
Chain-of-Thought Prompting
LLMs are known to respond more comprehensively by generating intermediate reasoning steps. In this section, we assess the effect of Zero-shot-CoT (Kojima et al., 2022) on VLind-Bench tasks by replacing the instruction in our prompts “Only respond in True or False.” to “Let’s think step by step.”.
Pipeline Score Accuracy
Models $S_{\textrm{CK}}$ $S_{\textrm{VP}}$ $S_{\textrm{CB}}$ $S_{\textrm{LP}}$ CB LP
Zero-shot-CoT
GPT-4o 91.4 94.7 93.1 89.4 93.0 87.8
GPT-4V 91.4 95.7 93.1 87.1 92.7 85.0
LLaVA-NEXT 72B 94.0 92.1 70.1 72.8 70.9 54.9
OmniLMM 12B 81.5 94.4 60.2 50.8 63.2 35.5
MiniCPM-V-2 2.8B 82.1 87.4 63.3 56.4 61.9 37.8
Table 5: Zero-shot-CoT performance of selected models. Compared to True/False prompting, improvements are shown in blue, while declines are shown in red.
As shown in Table 5, the impact of CoT varies depending on the model type and the scores measured. CoT produces significant performance improvements in certain large models, particularly in $S_{\textrm{LP}}$ and LP scores. However, in other instances, the advantages of CoT are negligible, and in some cases, CoT even hinders performance. Additionally, we found that smaller models, such as OminLMM and MiniCPM-V-2, struggled to effectively follow CoT instructions; they generated final answer before the reasoning steps. For these reasons, we adopted an evaluation setting that forces responses to be limited to either “True” or “False.”
@@
The tendency for the language prior to be inversely proportional to the scale of backbone LLMs may appear counterintuitive (i.e., LLaVA in Table 3). We have not identified the precise cause of this trend. One possible explanation is that larger pre-trained models are less prone to overfitting to the dataset during the visual instruction tuning process, thereby better maintaining their ability to attend to visual information.
In the experiments, we employ models with various scales of image encoders (ranging from approximately 300M to 5B), however, no clear correlation was observed between the language prior and the size of the image encoder.
@@
7 Conclusion
Limitations
Although VLind-Bench minimized potential confounding factors in assessing language priors, there may still be unconsidered factors that contribute to the benchmark scores. VLind-Bench used only a single fixed prompt for evaluation, but recent studies have shown that LLMs and LVLMs respond sensitively to even small changes in such prompts Zhou et al. (2024b); Lee et al. (2024).
Additionally, the CBs in our benchmark does not necessarily need to receive both text and image as input to check the commonsense bias. Such design choice is mostly due to a lack of established practices for feeding text-only inputs to LVLMs. As alternatives to $I_{\textrm{cf}}$, we conducted experiments using a plain single-color image or rendered text prompts as visual input (refer to Appendix B); however, none of these approaches works – these kinds of images can be considered out-of-distribution samples, and some proprietary models output error messages for these inputs. Exploring more established methods for text-only inputs in LVLMs falls outside the scope of our paper, but further research in this area is necessary both from a practical perspective and for a deeper understanding of how individual components of LVLMs operate.
Although VLind-Bench addressed various aspects of language priors and commonsense biases, its limitation is that it did not cover social bias Hall et al. (2023); Lee et al. (2023) or toxicity Kim et al. (2024); Koh et al. (2024).
Finally, we did not train the LVLMs with the data we constructed. While our primary goal in Section 4 was to generate data for a benchmarking purpose, we can also use this process to generate training data automatically. Training LVLMs with such dataset could help mitigate reliance on language priors, but we leave this as future work.
Acknowledgments
This work was partly supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (RS-2024-00348233), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)], and the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2024. K. Jung is with ASRI, Seoul National University, Korea.
@@ Appendix A Human Verification and Model Prompt Details
Appendix B Experiments Using a Plain White Image and Rendered Text Prompts
@@
LLaVA-NEXT 72B 94.4 95.7 76.1 58.6 75.5 46.7
OmniLMM 12B 88.1 97.7 78.6 81.4 79.5 66.4
@@
LLaVA-NEXT 72B 88.4 95.7 74.9 54.8 74.2 46.7
OmniLMM 12B 79.1 97.7 72.4 81.0 72.8 66.4
@@
LLaVA-NEXT 72B 89.1 95.7 70.6 54.2 72.8 46.7
OmniLMM 12B 74.2 97.7 65.2 77.4 68.2 66.4
@@
Appendix C Model Performance by Image Style
Here, we observed how performance varies across different image styles. As mentioned in Section 4, we generated images in photorealistic, illustration, and cartoon styles.
@@ Appendix C Model Performance by Image Style
LLaVA-NEXT 72B 94.6 95.8 77.2 65.0 76.5 52.4
OmniLMM 12B 88.8 97.7 81.8 82.8 83.1 70.5
@@
LLaVA-NEXT 72B 94.3 96.2 78.5 59.1 77.8 47.3
OmniLMM 12B 88.5 98.1 81.4 80.4 82.4 67.7
@@
LLaVA-NEXT 72B 94.8 95.5 78.8 58.2 78.8 48.0
OmniLMM 12B 87.7 97.8 82.2 82.0 82.5 68.4
Table 5 shows that the performance across these styles in the CK, VP, and CB did not vary significantly. A notable variation in performance was observed only in LP, where the photorealistic style yielded better results compared to the other two styles. This could be due to the model’s assessment that images in the illustration or cartoon styles lack realism compared to photorealistic images, leading it to generate responses that align more closely with common sense.
Appendix D Model Performance on Handcrafted Evaluation Set
To verify the validity of automatically generated text and images, we created a small handcrafted evaluation set and assessed the performance of several models, comparing it with their performance on the original VLind-Bench. All the text in the handcrafted evaluation set was written by humans, with three “triples” for each concept. For each triple, we generated three counterfactual images and one factual image. It was extremely challenging to find real images that depict counterfactual situations, and even if such images were found online, there was no way to verify whether they were outputs from generative models. To eliminate any potential advantage that OpenAI models might have from using DALL-E 3-generated images, we generated all the images using Stable Diffusion and Adobe Firefly, incorporating various styles such as photorealistic, illustration, and cartoon. This handcrafted evaluation set ultimately consists of 33 triples and 132 images, and the performance on this set is as follows.
Table 8: Experimental results on handcrafted evaluation set.
Pipeline Score Accuracy
Models $S_{\textrm{CK}}$ $S_{\textrm{VP}}$ $S_{\textrm{CB}}$ $S_{\textrm{LP}}$ CB LP
GPT-4o 90.9 90.9 100.0 93.8 100.0 93.9
GPT-4V 90.9 81.8 90.0 75.8 87.9 72.7
LLaVA-NEXT 72B 87.9 97.0 79.3 60.9 78.8 47.5
OmniLMM 12B 81.8 97.0 85.2 76.8 87.9 70.7
MiniCPM-V-2 2.8B 63.6 97.0 66.7 50.0 57.6 39.4
As shown in Table 8, the performance on the original VLind-Bench and the gold evaluation set does not differ significantly (refer to Table 3 for the original VLind-Bench scores).
Appendix E Data Samples
Here, we provide some data samples for each concept (next page). For the notations, please refer to the section 3.
Appendix F Information About Use Of AI Assistants
In writing this paper, we utilized ChatGPT  [^2] https://chatgpt.com/ for paraphrasing.